### v1

# MoE-DiffuSeq: Enhancing Long-Document Diffusion Models with Sparse Attention and Mixture of Experts

Abstract
In this paper, we present MoE-DiffuSeq, an innovative approach leveraging a Mixture of Experts (MoE) to enhance diffusion models for generating long documents. Traditional diffusion models, such as DiffuSeq, face significant computational inefficiencies and memory constraints when processing extended sequences. To overcome these challenges, we integrate a sparse attention mechanism with the MoE framework, facilitating efficient and scalable sequence generation. Our methodology incorporates a bespoke sparse attention framework designed to minimize computational complexity while preserving text quality and coherence. Additionally, we introduce a tailored soft absorbing state within the diffusion process to optimize the attention architecture, which accelerates sequence reconstruction and enhances precision. Empirical evaluations reveal that MoE-DiffuSeq substantially outperforms existing models in terms of training speed and sampling rates. These improvements are particularly beneficial for managing extended sequences, rendering our model apt for generating scientific articles, managing code repositories, and facilitating long dialogues. Benchmark comparisons further affirm that MoE-DiffuSeq enhances efficiency, speed, accuracy, and expressiveness, thereby advancing the capabilities of diffusion models for high-quality long-form text generation.
Intorduction
Recent advancements in neural network architectures, especially in the field of natural language processing (NLP), have significantly enhanced performance across a diverse array of tasks. The introduction of the Transformer model by vaswani2017attention; z1; z2 has revolutionized text processing by adeptly capturing contextual relationships within the text. However, traditional transformer architectures encounter substantial computational challenges when processing long sequences, primarily due to their inherent quadratic complexity. To address these challenges, architectures such as the Longformer (beltagy2020longformer; z3; z4) have been developed. These models incorporate sparse attention mechanisms, enabling more efficient management of extended sequences without compromising the model’s ability to handle complex dependencies.
Figure 1: Sparse Attention
Alongside advancements in transformer architectures, diffusion models have risen as a robust framework for generative tasks across multiple domains (chen-etal-2023-cheaper; bolliger-etal-2023-scandl; yuasa-etal-2023-multimodal; singh-etal-2023-codefusion; zhu-etal-2023-diffs2ut; zhu-etal-2023-diffusion; Z5; z6; z7). Originally developed for image synthesis, these models have been adeptly repurposed for sequence generation, exemplified by DiffuSeq (gong2023diffuseq). DiffuSeq harnesses diffusion principles to facilitate text generation, delivering competitive results while mitigating computational inefficiencies (gong2023diffuseq). Despite these advancements, significant challenges remain, notably in the areas of slow model convergence and the substantial computational demands of training diffusion models.
The challenge of efficiently handling long sequences in text generation is further accentuated by the inherent limitations of large language models (LLMs). Typically trained on text segments no longer than 8,000 tokens, these models often falter when tasked with processing significantly larger inputs, leading to a notable decline in performance (li2024long). Such limitations hinder their applicability in scenarios that demand a comprehensive understanding of Long Contexts, such as the transcription of scientific articles, generation of source code repositories, or management of extended dialogues. To mitigate these issues, various strategies have been employed, including the adoption of relative positional encodings and sliding-window attention mechanisms (beltagy2020longformer; li2023gaze; li2023internet; z8; z9). These approaches are designed to curtail memory consumption while preserving model effectiveness. However, they typically necessitate extensive retraining and may still fall short of fully addressing the fundamental challenges.
In this paper, we introduce MoE-DiffuSeq, a novel architecture that synergizes the Mixture of Experts (MoE) paradigm with the established DiffuSeq framework, augmented by a sparse attention mechanism (see in Figure 1). This fusion is designed to capitalize on the strengths of both systems, enhancing efficiency and scalability for generating lengthy documents. By integrating a bespoke sparse attention framework, MoE-DiffuSeq significantly reduces computational complexity while preserving the quality and coherence of the text. Furthermore, we incorporate a customized soft absorbing state within the diffusion process, which optimizes the attention architecture, thereby accelerating sequence reconstruction and enhancing precision. This methodological innovation not only streamlines the generative process but also ensures high fidelity in the output, making it a robust solution for extensive text generation tasks.
Our approach strategically addresses both the maintenance of high generative quality and the reduction of computational overhead, thereby optimizing performance for the efficient management of extensive textual inputs. By integrating sparse attention with the Mixture of Experts (MoE) framework, we significantly reduce the computational load. This reduction is achieved by minimizing the number of attention computations required, which is critical for optimizing the efficiency of iterative diffusion processes. This streamlined computation not only accelerates training convergence but also enhances sampling efficiency during inference, potentially surpassing the capabilities of traditional full-attention mechanisms in terms of generative quality. Such advancements establish our model as a robust solution for complex text generation tasks, ensuring both speed and accuracy without sacrificing the depth and coherence of the generated content.
Furthermore, MoE-DiffuSeq integrates cutting-edge advancements, including DPM-solver++ (lu2022dpm), to substantially reduce the number of diffusion steps required, thus accelerating the text generation process without sacrificing quality. This optimization proves particularly beneficial for applications such as machine translation and document summarization, where the demand for rapid, high-quality text generation is paramount. By enhancing the efficiency of the diffusion process, MoE-DiffuSeq enables more agile responses in dynamic environments, setting a new standard for performance in complex NLP tasks.
The principal contributions of this paper are summarized as follows:
  • Integration of Mixture of Experts (MoE) with DiffuSeq: We introduce MoE-DiffuSeq, a pioneering architecture that seamlessly integrates the Mixture of Experts paradigm into the DiffuSeq framework. This synthesis significantly enhances the efficiency and scalability of diffusion models tailored for generating extensive documents. By dynamically selecting the most relevant experts for varying segments of the text, our model adeptly manages the computational complexities traditionally associated with processing long sequences.
  • Incorporation of Sparse Attention Mechanism: Our methodology incorporates a tailored sparse attention mechanism within the MoE-DiffuSeq model, significantly reducing the number of attention computations needed. This reduction not only lowers the overall computational burden but also ensures the efficient processing of extensive textual inputs without sacrificing the quality or coherence of the generated text, which is essential for producing high-quality long documents.
  • Empirical Validation and Performance Improvements: Through rigorous empirical evaluations, MoE-DiffuSeq has demonstrated superior performance over existing methods across various NLP tasks, particularly in handling long texts with reduced computational demands. The integration of state-of-the-art techniques, such as DPM-solver++, further augments our model’s efficiency by minimizing the necessary diffusion steps, thereby speeding up both the training and sampling phases without compromising the quality of text generation.
These contributions underline our model’s innovative approach to combining robust machine learning paradigms and optimizing them for advanced text generation tasks, setting a new benchmark for future research in the field.
Related Work
In the rapidly evolving field of text generation, diffusion models have become prominent due to their unique ability to handle complex generative tasks. This section provides an overview of the related work, placing our contributions within the broader context of recent advancements in the field.
Mixture of Experts
The Mixture of Experts (MoE) approach has revolutionized the scalability of neural networks by dynamically allocating computational resources across a diverse set of expert networks (gao2022parameter; du2022glam; z10). This paradigm, initially introduced by rajbhandari2022deepspeed; z11; z12, employs a gating mechanism that selectively activates the most relevant experts for a given input, significantly enhancing computational efficiency and model performance, particularly in large-scale applications.
Recent enhancements in MoE technology, such as the Switch Transformer (fedus2022switch), have demonstrated its capability to handle extensive datasets with substantially reduced computational overhead. The Switch Transformer, for instance, has proven effective in training models with trillions of parameters by activating only a subset of the model’s parameters for each input. Furthermore, the GShard framework (lepikhin2020gshard; z13; z14) advances MoE’s potential by enabling the efficient training of very large models through a strategic combination of expert routing and sharding. This innovation underlines the suitability of MoE for managing the complexity and scalability challenges in various NLP tasks (zhou2022mixture; z15; z16).
Building on these developments, our work integrates MoE with the DiffuSeq framework, aiming to enhance the efficiency and scalability of diffusion models tailored for extensive document generation. By dynamically selecting relevant experts for different text segments, our model effectively addresses the computational complexity associated with processing long sequences. This ensures that the generative process remains efficient and scalable, even as sequence lengths increase, making it particularly advantageous for producing detailed scientific documents, extensive code repositories, and comprehensive narratives.
Sparse Attention Mechanisms
Sparse attention mechanisms have become a cornerstone in optimizing transformer architectures for efficient processing of long text sequences. The Longformer architecture (beltagy2020longformer; z17; z18) represents a significant evolution in this field. It combines local windowed attention with strategically placed global attention mechanisms, effectively managing extended documents. This hybrid attention model drastically reduces the inherent computational complexity that scales quadratically with sequence length in conventional full attention mechanisms. Local attention focuses on processing nearby tokens efficiently, while global attention spans the entire sequence, maintaining a holistic understanding of the context. This dual approach is particularly vital for comprehensive tasks such as document summarization and extensive question answering.
Other innovative sparse attention models include BigBird (zaheer2020big) and ETC (ainslie-etal-2020-etc), each enhancing performance for specific NLP tasks through unique attention schemes. BigBird extends the sparse attention concept by integrating random and global attention with local attention, facilitating the handling of even longer sequences with reduced computational demands. Conversely, ETC optimizes the processing of structured data by melding local and global attention mechanisms within a hierarchical model structure, further refining efficiency and scalability.
Our methodology integrates these advanced sparse attention configurations within a diffusion-based framework for sequence generation, specifically addressing computational efficiency and enhancing the fidelity of the generated sequences. By adopting sparse attention, our model processes extensive documents more effectively, ensuring optimal utilization of computational resources without compromising the quality of the output. This integration not only alleviates the computational burden but also bolsters the model’s capability to produce coherent and contextually accurate long-form text, setting a new standard in the field of text generation.
Diffusion Models for Text Generation
Diffusion models have rapidly emerged as a formidable alternative to traditional generative models, effectively modeling text in continuous latent spaces. Central to our study is the DiffuSeq framework (gong2023diffuseq; z19), which exemplifies the sophisticated application of diffusion models tailored for text generation. These models refine a noisy initial input iteratively, gradually transforming it into coherent and structured text. This process facilitates the creation of high-quality content that adeptly captures complex dependencies and structures inherent within textual data.
Historically, pioneering works such as those by hoogeboom2021autoregressive; z20 and austin2021structured have expanded the utility of diffusion models. hoogeboom2021autoregressive; z21 explored character-level text generation using an autoregressive diffusion model approach, showcasing the potential for capturing nuanced textual details. Similarly, austin2021structured developed structured denoising diffusion models that incorporate a novel absorbing state concept, significantly aiding in maintaining the coherence and structural integrity of the generated text.
Despite the innovative strides in this domain, earlier models frequently encountered obstacles when addressing longer or more complex textual sequences, primarily due to the limitations imposed by their foundational designs, which were often discrete or overly simplistic. To overcome these challenges, our work integrates sparse attention mechanisms within the DiffuSeq framework. This integration not only leverages the inherent strengths of diffusion models in generating high-quality text but also introduces a refined method to manage the increased demands of sequence length and complexity effectively. Through this synergy, our model, MoE-DiffuSeq, is equipped to handle extended sequences more efficiently, ensuring the generation of text that is coherent and contextually appropriate across broader narratives. This advancement significantly enhances the model’s applicability in generating detailed and expansive documents, setting a new benchmark in the field of text generation.
Method
In this section, we detail the methodology employed to enhance DiffuSeq for long document generation by leveraging a Mixture of Experts (MoE) framework and incorporating a sparse attention mechanism. The proposed model, MoE-DiffuSeq (in Figure 2), is designed to efficiently handle the increased complexity and length of textual data while maintaining high performance.
Figure 2: Model Architecture
Integration of Sparse Attention Mechanism
To address the computational complexities associated with generating extended documents, our model, MoE-DiffuSeq, incorporates a sparse attention mechanism inspired by the Longformer architecture. Traditional Transformer models are well-known for their quadratic computational expense ($O(n^{2})$) relative to the sequence length $n$. This becomes impractical for long sequences due to escalating computational demands and memory constraints.
To mitigate these issues, we implement a sliding window attention mechanism, where each token computes attention weights only for a constrained set of neighboring tokens within a predefined window. This approach effectively reduces the computational complexity to $O(n\times w)$, where $w$ denotes the window size, making the processing of long sequences computationally feasible.
The attention computation within our sparse mechanism is defined as follows:
$\text{Attention}(Q_{i},K_{j},V_{j})=\text{softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right)V_{j},$
where: - $Q_{i}$, $K_{j}$, and $V_{j}$ represent the query, key, and value vectors for tokens $i$ and $j$, respectively. - $d_{k}$ is the dimensionality of the key vectors.
This formulation ensures that each query vector $Q_{i}$ interacts only with key vectors $K_{j}$ within its designated window, significantly streamlining the computational load by focusing processing power where it’s most needed.
Dilated Sliding Window for Long Context
To effectively capture long-range dependencies within extensive text sequences, we have implemented a dilated sliding window mechanism in our MoE-DiffuSeq architecture. This approach does not increase the size of the attention window $w$, yet it expands the effective receptive field by introducing gaps, denoted by $d$, between tokens within the window. This method allows each attention head to process a broader context without a corresponding increase in computational load.
The extended receptive field facilitated by the dilated sliding window is mathematically represented as follows:
$\text{Receptive Field}=l\times d\times w,$
where:
  • $l$ represents the number of transformer layers,
  • $d$ is the dilation factor, determining the gap between tokens in the attention mechanism,
  • $w$ denotes the window size, specifying the number of tokens each attention head can directly interact with.
This configuration allows the model to integrate information over larger textual spans effectively, thus enhancing its ability to understand and generate coherent long-form content. The dilated window approach ensures that critical linguistic structures, which may be spread out across large segments of text, are captured and processed efficiently.
Global Attention for Key Tokens
MoE-DiffuSeq also employs global attention to specific key tokens, such as the [CLS] token for classification tasks or particular tokens in question-answering contexts. This ensures that these tokens can attend to all other tokens in the sequence, integrating comprehensive contextual information. The global attention mechanism is defined as:
$\text{GlobalAttention}(Q_{g},K_{g},V_{g})=\text{softmax}\left(\frac{Q_{g}K_{g}^{T}}{\sqrt{d_{k}}}\right)V_{g},$
where $Q_{g}$, $K_{g}$, and $V_{g}$ are the query, key, and value matrices for tokens designated for global attention.
Incorporation of Mixture of Experts
To further enhance model capacity and efficiency, we incorporate a Mixture of Experts (MoE) framework. Each layer in the transformer is augmented with multiple expert networks, and a gating mechanism dynamically selects a subset of these experts for each input token. The gating function $G(x)$ determines the probability $p_{i}$ of selecting expert $i$ based on the input token $x$:
$G(x)=\text{softmax}(W_{g}x),$
where $W_{g}$ is the gating weight matrix. The output of the MoE layer is a weighted sum of the selected experts’ outputs:
$\text{MoE}(x)=\sum_{i}p_{i}E_{i}(x),$
where $E_{i}$ represents the $i$-th expert.
Adaptation of Diffusion Processes
We adapt the diffusion process from DiffuSeq to incorporate Gaussian noise and a discrete absorbing state. The forward diffusion process is modified as:
$z_{t}=\sqrt{\overline{\alpha}_{t}}z_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,$
where $\epsilon$ denotes Gaussian noise, and $z_{t}$ is the state at time $t$. A discrete absorbing state $m$ is introduced probabilistically, enhancing the model’s ability to manage the granularity of textual data.
Joint Denoising and Loss Optimization
During the reverse process, a joint denoising strategy reconstructs the data from both continuous and discrete noise. The loss function for this process is formulated as:
$\mathcal{L}=\sum_{t=2}^{T}\|\text{EMB}(w_{t})-f_{\theta}(z_{t},t)\|^{2}+R(\|z_{0}\|),$
where EMB is the embedding function converting discrete tokens into continuous embeddings, and $f_{\theta}$ is the denoising function. The regularization term $R(\|z_{0}\|)$ ensures stability and quality in the generative process.
Consistency in Sampling and Inference
To ensure consistency between training and inference, the same noise model is used in both phases. The reverse diffusion process during sampling is calculated through an integral formulation:
$z_{t}=z_{s}+\int_{s}^{t}e^{f_{\theta}(z,d)}\,dd,$
where $z_{s}$ represents the state at time $s$, and $e^{f_{\theta}(z,d)}$ denotes the exponentiated output of the denoising function. This integral is evaluated using the Euler method, ensuring numerical stability and accurate sequence reconstruction.
Computational Efficiency and Scalability
The integration of sparse attention and the MoE framework significantly enhances computational efficiency, especially for longer sequences. This combination allows MoE-DiffuSeq to scale effectively across diverse datasets and varying sequence lengths, setting a new standard for efficiency and scalability in generative modeling for NLP tasks.
Experiments
Experimental Setup
To rigorously evaluate the performance of MoE-DiffuSeq in generating long documents, we employed four diverse datasets, each chosen for its unique challenges in natural language generation. The Arxiv Dataset (cohan-etal-2018-discourse) allowed us to assess the model’s ability to generate coherent and structured scientific documents. In contrast, the HotpotQA dataset (yang-etal-2018-hotpotqa) tested the model’s capacity to maintain contextual integrity and reasoning across extended interactions. The Commonsense Conversation Dataset (zhou-etal-2021-commonsense) provided a platform to evaluate the generation of contextually appropriate and pragmatic dialogue responses. Lastly, the QQP dataset (wang2017bilateral) measured the model’s paraphrasing capabilities, focusing on its ability to retain semantic meaning while altering phrasing.
Each dataset necessitated specific evaluation metrics tailored to measure the model’s performance against the distinct challenges posed by the dataset. This structured approach allowed for both quantitative and qualitative analysis of the model’s capabilities, ensuring a comprehensive assessment of MoE-DiffuSeq’s effectiveness in handling the complexities of generating long-form text across various domains.
Baselines and Comparative Analysis
To rigorously assess the effectiveness of MoE-DiffuSeq in long document generation, we conducted a comparative analysis with several leading models renowned for their text-generation capabilities:
  • DiffuSeq (gong2023diffuseq): Serving as the foundational architecture, DiffuSeq offers a direct baseline, allowing us to underscore the enhancements brought about by integrating the Mixture of Experts (MoE) framework and sparse attention mechanisms.
  • Longformer (beltagy2020longformer): Recognized for its adept handling of extensive texts via sparse attention, Longformer provides a benchmark to gauge the incremental benefits our MoE-DiffuSeq introduces, particularly in managing extensive sequence lengths efficiently.
  • GPT-4 (achiam2023gpt): As a benchmark in generative tasks, GPT-4 helps establish a high-performance standard, showcasing MoE-DiffuSeq’s competitive stance in the landscape of advanced text generation technologies.
In addition to these comparisons, MoE-DiffuSeq’s performance was meticulously evaluated against specialized models tailored for each specific dataset used in our study. This approach not only highlights MoE-DiffuSeq’s adaptability across various natural language processing challenges but also provides a transparent view of its performance nuances in distinct task environments. This comprehensive evaluation strategy ensures a well-rounded analysis of MoE-DiffuSeq’s capabilities and advancements in text generation.
Implementation Details
MoE-DiffuSeq integrates 12 Transformer layers with 12 attention heads per layer, utilizing Longformer’s sparse attention mechanism within the DiffuSeq framework. The model employs a Mixture of Experts (MoE) approach, where each layer includes multiple expert networks, and a gating mechanism dynamically selects the most relevant experts for each input token. This gating function $G(x)$ is defined as:
$G(x)=\text{softmax}(W_{g}x),$
where $W_{g}$ is the gating weight matrix. The output of the MoE layer is a weighted sum of the selected experts’ outputs:
$\text{MoE}(x)=\sum_{i}p_{i}E_{i}(x),$
where $p_{i}$ is the probability of selecting expert $i$ and $E_{i}$ denotes the $i$-th expert.
The training was conducted using a staged approach, gradually increasing window sizes and sequence lengths. We utilized 2,048 diffusion steps with a square-root noise schedule, optimizing the balance between computational efficiency and text generation quality. The forward diffusion process is represented by:
$z_{t}=\sqrt{\overline{\alpha}_{t}}z_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,$
where $\epsilon$ denotes Gaussian noise, and $z_{t}$ is the state at time $t$.
Evaluation Metrics
To rigorously evaluate the performance of MoE-DiffuSeq and baseline models, we employed a comprehensive set of evaluation metrics. These metrics are designed to assess different dimensions of text generation quality, including linguistic coherence, diversity, and semantic accuracy:
  • BLEU (Bilingual Evaluation Understudy) (papineni2002bleu): Measures the phrase-level accuracy between machine-generated text and human-written references, providing insights into the precision of the generated text.
  • ROUGE (Recall-Oriented Understudy for Gisting Evaluation) (lin2004rouge): Evaluates the overlap of n-grams between the generated text and the reference text, with a focus on recall, to assess the completeness of the generated content.
  • BERTScore (zhang2019bertscore): Utilizes BERT’s pre-trained contextual embeddings to measure the cosine similarity between words in the candidate and reference sentences, offering a measure of semantic similarity that reflects the contextual accuracy of the generated text.
These metrics collectively provide a robust framework for evaluating various dimensions of text generation quality, including linguistic coherence, diversity, and semantic accuracy.
Data Handling and Analysis
For each dataset, we generated multiple text samples per input using MoE-DiffuSeq and the baseline models. This allowed us to compute diversity metrics, assessing the variety and richness of the generated text. Experiments were conducted on NVIDIA A100 GPUs to ensure optimal performance and fair comparison across models.
The results were analyzed to determine the model’s ability to generate high-quality, coherent, and contextually appropriate long-form text. We observed that MoE-DiffuSeq consistently outperformed the baselines in maintaining long-range dependencies and generating text with higher semantic accuracy and diversity.
Results and Analysis
In this section, we present the results and analysis of our experiments using the MoE-DiffuSeq model, which incorporates the Mixture of Experts (MoE) framework and a sparse attention mechanism to enhance long document generation. We evaluated our model on several datasets, including the Arxiv Abstract Dataset, HotpotQA, Commonsense Conversation Dataset, and Quora Question Pairs (QQP). The primary evaluation metrics were BLEU, ROUGE, and BERTScore.
Main Results
Our experiments demonstrate that the MoE-DiffuSeq model consistently outperforms previous models, including the Longformer and DiffuSeq, across various datasets. The following tables provide a detailed comparison of the performance metrics.
Arxiv Abstract Dataset
Based on the experimental results, MoE-DiffuSeq demonstrates superior performance in generating coherent and contextually accurate summaries of scientific texts on the Arxiv Abstract Dataset. As shown in Table 1, MoE-DiffuSeq achieves the highest scores across all metrics (R1: 44.41, R2: 18.73, RL: 39.89), outperforming both Longformer and DiffuSeq. This indicates its robust capability to handle the complexities of scientific language and structure, making it an excellent choice for summarizing scientific literature.
Model R1 R2 RL
Longformer 41.44 17.52 38.70
DiffuSeq 39.12 16.43 37.88
MoE-DiffuSeq 44.41 18.73 39.89
Table 1: Performance comparison on the Arxiv Abstract Dataset.
HotpotQA Dataset
On the HotpotQA dataset, MoE-DiffuSeq exhibits substantial improvements in both Answer EM/F1 and Support EM/F1 scores, as shown in Table 2. MoE-DiffuSeq achieves an Answer EM/F1 of 72.88 / 85.42 and a Support EM/F1 of 66.69 / 90.40, outperforming Longformer and DiffuSeq. These results underscore MoE-DiffuSeq’s robustness and effectiveness in handling complex, multi-hop question-answering tasks, making it a promising model for applications requiring nuanced understanding and synthesis of information across multiple documents.
Model Answer EM/F1 Support EM/F1
Longformer 71.21 / 82.42 65.11 / 89.50
DiffuSeq 70.91 / 81.43 64.60 / 88.51
MoE-DiffuSeq 72.88 / 85.42 66.69 / 90.40
Table 2: Performance comparison on the HotpotQA Dataset.
Commonsense Conversation Dataset
For the Commonsense Conversation Dataset, MoE-DiffuSeq achieves superior performance across BLEU, ROUGE-L, and BERTScore metrics, as depicted in Table 3. With scores of 0.049 for BLEU, 0.233 for ROUGE-L, and 0.628 for BERTScore, MoE-DiffuSeq outperforms both Longformer and DiffuSeq. This indicates its effectiveness in generating diverse and contextually appropriate conversational responses, highlighting its potential for applications in dialogue systems and conversational AI.
Model BLEU ROUGE-L BERTScore
Longformer 0.030 0.139 0.602
DiffuSeq 0.022 0.119 0.501
MoE-DiffuSeq 0.049 0.233 0.628
Table 3: Performance comparison on the Commonsense Conversation Dataset.
Quora Question Pairs (QQP)
Model Accuracy
Longformer 92.3
DiffuSeq 91.7
MoE-DiffuSeq 95.3
Table 4: Accuracy comparison on the QQP Dataset.
In the QQP dataset, MoE-DiffuSeq outperforms other models in terms of accuracy, as illustrated in Table 4. Achieving an accuracy of 95.3, MoE-DiffuSeq demonstrates its superior paraphrasing capabilities compared to Longformer and DiffuSeq. This highlights its effectiveness in generating precise and accurate paraphrases, making it highly suitable for tasks requiring nuanced understanding and rephrasing of text.
Comparative Discussion
The integration of sparse attention with the diffusion model in MoE-DiffuSeq has significantly enhanced its performance across all tested datasets. The model not only excels in handling longer sequences but also shows marked improvements in metrics evaluating semantic coherence and factual accuracy. These results support our hypothesis that the hybrid approach, leveraging the strengths of both sparse attention and diffusion models, provides superior performance in complex NLP tasks.
Ablation Study
To understand the contributions of individual components within the MoE-DiffuSeq model, we conducted an ablation study by modifying the sparse attention component, the number of diffusion steps, and the attention window sizes. The baseline MoE-DiffuSeq model combines sparse attention with DiffuSeq. As shown in Table 5, the study includes various configurations such as removing sparse attention, altering the diffusion steps, and changing the attention window sizes to evaluate their impact on performance metrics like BLEU, ROUGE, and BERTScore.
Configuration Attention Type Diffusion Steps Window Size BLEU/ROUGE/BERTScore
Baseline (Full Model) Sparse 2048 512 44.41/18.73/39.89
No Sparse Attention Standard 2048 512 42.52/17.99/38.41
Reduced Diffusion Steps Sparse 1024 512 43.11/18.03/39.26
Increased Diffusion Steps Sparse 4096 512 44.71/18.55/40.20
Smaller Window Size Sparse 2048 256 43.80/18.22/39.65
Larger Window Size Sparse 2048 1024 44.40/18.66/39.92
Table 5: Ablation study results comparing different configurations of the MoE-DiffuSeq model on the Arxiv dataset.
These results underscore the critical balance between attention mechanisms and diffusion steps for optimal performance. Sparse attention is essential for effectively handling long sequences, as evidenced by the significant drop in performance when it is removed. Adjustments to the diffusion steps indicate that more steps can enhance text coherence, but beyond a certain point, the improvements are marginal. Similarly, larger attention windows provide a broader contextual range, slightly improving performance, though the benefits are less pronounced than those from incorporating sparse attention. This highlights the importance of carefully tuning these components to maximize the model’s effectiveness in generating coherent and contextually accurate summaries of scientific texts.
Comparative Analysis on Arxiv Dataset
Figure 3: Comparison of Model Performance on Inference and Novelty
In the evaluation of 2-gram novelty and inference time on the Arxiv dataset (Figure 3), notable differences emerge among Longformer and DiffuSeq. MoE-DiffuSeq demonstrates a superior balance between high novelty in generated text and efficient inference times. Specifically, MoE-DiffuSeq maintains a higher 2-gram novelty score compared to its competitors, suggesting it generates more unique and varied bi-grams, crucial for producing diverse and innovative textual outputs. Despite its high novelty score, MoE-DiffuSeq’s inference time remains competitive, only slightly slower than Longformer, which boasts the fastest inference but at the cost of significantly lower novelty.
Overall, DiffuSeq shows the lowest performance in both metrics, indicating potential areas for improvement in its model architecture or optimization processes. Comparatively, Longformer, while excelling in speed, falls behind in generating novel text sequences, limiting its utility in applications requiring high creativity and variation in text output.
Conclusion
In this study, we introduce MoE-DiffuSeq a novel model that integrates a Mixture of Experts (MoE) framework with DiffuSeq, specifically designed to enhance long document generation through sparse attention mechanisms. Our approach effectively addresses the computational challenges associated with long-sequence text generation while improving the overall quality of the generated content. Through comprehensive experiments on multiple datasets, MoE-DiffuSeq has demonstrated notable improvements in both efficiency and output quality compared to existing models. These promising results underscore the potential of combining sparse attention with diffusion models for advanced text generation tasks. Future work will aim to expand the applicability of MoE-DiffuSeq to other domains and explore the incorporation of multimodal inputs to further extend its utility. The encouraging outcomes of this study pave the way for continued exploration and refinement of hybrid generative models leveraging sparse attention and diffusion techniques.
Appendix A Preliminaries
Diffusion Models for Text Generation
Diffusion models have emerged as a robust framework for generative processes, particularly in the domain of text generation, where models like DiffuSeq exemplify their potential. These models operate through a two-phase approach: the forward diffusion process and the reverse diffusion process.
Forward Diffusion Process
During the forward diffusion phase, an initial data point, $x_{0}$, undergoes a transformation into a noise-dominated state through a sequence of intermediate states $x_{t}$ for $t=1,2,\ldots,T$, where $T$ signifies the total number of timesteps. The mathematical representation of this process is given by:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
where $\beta_{t}$ denotes the variance schedule, typically a small, incrementally increasing positive value. As $t$ increases, $x_{t}$ progressively approaches a purely noisy state, ultimately reaching $x_{T}\sim\mathcal{N}(0,I)$, symbolizing a standard Gaussian distribution.
Reverse Diffusion Process
The objective of the reverse diffusion process is to reconstruct the original data from its noisy counterpart. This reconstruction is facilitated by a learned denoising function $f_{\theta}(x_{t},t)$, which predicts $x_{t-1}$ from $x_{t}$:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};f_{\theta}(x_{t},t),\sigma_{t}^{2}I),$
where $\sigma_{t}$ represents a learned variance parameter. The process iteratively refines $x_{T}$, moving towards $x_{0}$ to effectively reverse the noise addition.
Handling Discrete States
In scenarios where the states are discrete, such as text tokens, each state $x_{t}$ is treated as a distinct entity. The forward diffusion might employ a categorical distribution, while the reverse process may involve transitions between discrete states, often facilitated by methods like masking or token substitution.
This dual-phase diffusion approach not only allows for the effective generation of high-quality text that captures intricate dependencies but also adapts well to the inherent complexities of structured and unstructured data. By leveraging the diffusion model framework within the DiffuSeq architecture, we can address significant challenges in text generation, particularly those associated with the management of long sequences and the preservation of semantic integrity.
Mixture of Experts (MoE) in Text Generation
The Mixture of Experts (MoE) paradigm revolutionizes large-scale neural network architectures by dynamically allocating computational resources among specialized sub-models, or ”experts,” to process each input. This modular approach significantly enhances scalability and adaptability, particularly when managing diverse and complex tasks.
Mathematical Representation
An MoE model is mathematically represented as:
$y=\sum_{i=1}^{N}g_{i}(x)E_{i}(x),$
where $N$ denotes the total number of experts, $g_{i}(x)$ is the gating function that allocates weights to the $i$-th expert $E_{i}$, based on the input $x$. The gating function is designed to produce a sparse output, ensuring that only a relevant subset of experts is activated for any given input, thereby optimizing computational resources.
Gating Mechanism
The gating mechanism, pivotal for the efficiency of MoE, is expressed as:
$g_{i}(x)=\frac{\exp(h_{i}(x))}{\sum_{j=1}^{N}\exp(h_{j}(x))},$
where $h_{i}(x)$ is a learnable scoring function that evaluates the suitability of each expert for the current input. This mechanism promotes sparsity and specialization by activating only the most relevant experts, thus reducing unnecessary computations.
Application in Long Text Generation
In the context of long text generation, MoE models are particularly advantageous. They dynamically allocate processing power to handle various segments of a text, adapting to the unique demands of each segment. This ability is crucial for generating coherent and contextually accurate extended texts, such as detailed scientific articles, elaborate narratives, and comprehensive reports.
By integrating MoE into the DiffuSeq framework, our approach—MoE-DiffuSeq—not only leverages the inherent strengths of diffusion models but also introduces a scalable method to manage the increased computational demands of processing lengthy documents. This integration ensures that our model can efficiently handle sequences of considerable length without sacrificing the quality of the generated content.
This dynamic allocation of computational resources, tailored to the specific requirements of different text segments, ensures that our generative process remains both efficient and effective, making MoE-DiffuSeq a cutting-edge solution in the realm of advanced text generation systems.
Sparse Attention Mechanism
Efficiently managing long text sequences presents a significant challenge in natural language processing. The sparse attention mechanism, exemplified by architectures like Longformer, directly addresses this by reducing the computational overhead traditionally associated with the self-attention mechanism of Transformer models. This reduction is achieved by altering the complexity from $O(n^{2})$ to $O(n\times w)$, where $n$ represents the sequence length and $w$ denotes the window size.
Mathematical Formulation
The sparse attention mechanism is mathematically defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
where $Q$, $K$, and $V$ represent the query, key, and value matrices respectively, and $d_{k}$ is the dimensionality of the key vectors. In this configuration, each token computes attention weights only for a subset of tokens within a predefined window, significantly reducing the number of operations involved.
Implementation of Sparse Attention
Sparse attention is implemented as:
$\text{Sparse Attention}(Q,K,V)=\sum_{i=1}^{n}\sum_{j\in\mathcal{W}(i)}\text{softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right)V_{j},$
where $\mathcal{W}(i)$ specifies the set of indices within the attention window of the token $i$. This method efficiently limits the scope of attention, focusing computational resources where they are most needed.
Enhancements for Long Context
To capture dependencies over longer distances without increasing computational demands, dilated attention is utilized:
$\mathcal{W}(i)=\{i-k\cdot d,\ldots,i+k\cdot d\},$
where $d$ represents the dilation factor, enhancing the model’s ability to integrate information over larger textual spans without a corresponding increase in the window size.
Additionally, global attention tokens are incorporated to provide context integration across the entire sequence. This feature is particularly valuable in tasks that require an overarching understanding of the full text, such as in document summarization or complex question-answering scenarios.
Advantages of Sparse Attention
Integrating sparse attention mechanisms allows models to efficiently process extensive documents while ensuring the preservation of text quality and coherence. This approach not only mitigates the quadratic scaling issue but also enhances the model’s capability to generate contextually rich and coherent long-form text. By focusing computational resources judiciously, sparse attention mechanisms facilitate deeper and more effective text analysis and generation, especially in applications involving substantial textual data.